A fence open on both sides has 26 wooden posts each 10 feet apart. What is the length of the fence?
Found 3 solutions by ikleyn, Alan3354, greenestamps:
Answer by ikleyn(52781)

.
A fence open on both sides has 26 wooden posts each 10 feet apart.
What is the length of the fence?
~~~~~~~~~~~~~~~~~

After reading the post, I can assume that the fence is along the entire perimeter
of 10*26 = 260 feet, except of two open sections/parts between the wooden posts.


If so, then the length of the fence is  10*26 - 2*10 = 260 - 20 = 240 feet.    ANSWER

Solved.




Answer by Alan3354(69443)

"A fence open on both sides..."
------------
What does that mean?
Are some fences closed?

Answer by greenestamps(13200)



As indicated in the responses from the other two tutors, the wording of the problem is abysmal. Re-post the problem exactly as it was given to you; if this is the way it was given to you, then there's not much we (or you) can do, except guess the real meaning.

"A fence open on both sides" could be simply a straight length of fence. In that case, 26 posts define 25 lengths of fence. So perhaps the length of the fence is 25*10 = 250 feet?

Note, however, that the phrase "each 10 feet apart" is terrible language also. It almost certainly means that the distance between neighboring posts is 10 feet; but that's not what the words say. "apart" indicates that you are talking about two (or more) objects; grammatically, "each (one) 10 feet apart" is meaningless.

ANSWER (my guess): 250 feet